How many grams of fat are needed to make up 10% of a 900 Calorie meal
Answer #1

10% of the 900 calories = 900*10/100= 90 calories

Since 9 calories of energy is provided by 1gram of fat

90 calories will be provided by = (1/9)*90 gram = 10 grams of fat     Answer